OpenStudy (anonymous):

wat is 39x60 plzz show work

myininaya (myininaya):

what is 39*6

OpenStudy (anonymous):

you can rewrite it as 390 * 6= 2340

OpenStudy (anonymous):

ya but thats harder not wat im looking 4

OpenStudy (anonymous):

39 = 3*13 60 = 6*10 = 3*2*5*2 \[\implies 39 \times 60 = 3^2*2^2 * 5 * 13\] Take your pick on what order you want to do em in.

OpenStudy (anonymous):

no its easier now you have 390 6 * (18+5) 40 2340

OpenStudy (anonymous):

(40-1)*60 2400-60 2340

myininaya (myininaya):

oops i forgot to mark out the 2nd row for the second on i did

myininaya (myininaya):

another example coming up

myininaya (myininaya):

there are more ways fun ways to multiply but i forgot them when i finished math history